# smile: Single-turn to Multi-turn Inclusive Language Expansion via ChatGPT for Mental Health Support

### v2

@@ Abstract
Developing specialized dialogue systems for mental health support requires multi-turn conversation data, which has recently garnered increasing attention. However, gathering and releasing large-scale and real-life multi-turn conversations to facilitate advancements in mental health presents challenges due to data privacy protection, as well as the time and cost involved. To address the challenges related to data scarcity, we introduce smile, a single-turn to multi-turn inclusive language expansion technique that prompts ChatGPT to rewrite public single-turn dialogues into multi-turn ones. Our work begins with the analysis of language transformation, validating the feasibility of the proposed method when compared with other baseline methods. We then conduct a study on dialogue diversity, including lexical features, semantic features, and dialogue topics, demonstrating the effectiveness of our proposed method. Furthermore, we implement an expert evaluation and the results demonstrate that the dialogues generated with our proposed method are of higher quality than those generated with other baseline methods. Thus, we employ our method to generate a large-scale, diverse, and high-quality dialogue dataset named SmileChat, comprising 55,165 dialogues in total with an average of 10.4 turns per dialogue. Finally, we utilize the collected corpus to develop a mental health chatbot, MeChat. To better assess the overall quality of SmileChat, we collect a real-life chat dataset comprising 82 counseling dialogues for model evaluation. Both automatic and human evaluations demonstrate that our trained dialogue system exhibits significant improvements, showcasing that SmileChat is high-quality and practical.
@@ 1 Introduction
Figure 1: Example of instructing ChatGPT to rewrite a single-turn dialogue into a multi-turn one.
@@
We all know the importance of mental health and mental health issues (Kessler et al., 2005) have been a persistent concern for human beings. Recently, advancements in natural language processing (NLP) technology (Vaswani et al., 2017; Ouyang et al., 2022; Ni et al., 2022) have led to the emergence of neural-based conversational AI applied in various domains, including mental health (Liu et al., 2022; Tu et al., 2022). Virtual counselors powered by AI, as an innovative solution for mental health, can effectively address accessibility barriers, such as the high cost of treatment and the shortage of trained professionals to meet the demand. Furthermore, such a dialogue system provides mental health support as an effective and practical online counseling approach for those in need, safeguarding user privacy and mitigating the stigma that often accompanies help-seeking. However, the lack of publicly available, large-scale, diverse, and high-quality multi-turn chat datasets in the mental health support domain hinders the development of specialized dialogue systems.
@@
Motivation Conversations related to mental health support often contain sensitive information and must be kept confidential (Lu et al., 2021) to safeguard the privacy of individuals seeking help. Making these conversations publicly available may discourage individuals from seeking support or negatively impact their personal and professional lives once known to people they are acquainted with. To facilitate progress in the NLP community, some researchers have attempted to collect various dialogue corpora (Liu et al., 2021; Sun et al., 2021; Zheng et al., 2022) through crowd-sourcing, data crawling, or data augmentation to build a dialogue agent capable of providing emotional and mental health support. How to construct a large-scale, diverse and high-quality multi-turn chat dataset for mental health motivates us to carry out the work as presented in this paper.
@@
Challenges To be more specific, crowd-sourcing conversations (Liu et al., 2021) for emotional support has limitations due to the high cost and time required to train and manage annotators, as well as the difficulty in mimicking real-life interactions, that is, interlocutors may lack an understanding of the dilemma of living with mental disorders. An alternative is crawling QA (Sun et al., 2021) on a public mental health forum for training psychological support models. However, single-turn conversations may not be sufficient for resolving mental health issues, as multiple interaction exchanges are often needed. Multi-turn conversations, which can better simulate real-world conversations, are therefore more practical for training psychological support models. While the post-triggered machine-augmented method (Zheng et al., 2022) can address the limitations of scale and topic diversity, it does not take into account the responses of experienced supporters.
@@
Our Approach To tackle the challenges mentioned above, we introduce smile, single-turn to multi-turn inclusive language expansion via ChatGPT. Specifically, we instruct ChatGPT to transform publicly available question-answer pairs (public QA), which can also be considered as single-turn dialogues from the real world, into multi-turn conversations, as illustrated in Figure 1. With the proposed method, we build a large-scale, diverse, and high-quality multi-turn conversation dataset for mental health support.
@@
[…]
Our Contributions We make our data, code, and model publicly available. We believe our work offers a new perspective on constructing a large-scale, diverse, and high-quality multi-turn dialogue dataset for mental health within the research community. Our contributions can be summarized as follows:
  • We introduce smile, which provides a novel perspective for alleviating the scarcity of multi-turn conversations in mental health.
  • Through the analysis of language transformation and dialogue diversity, we verify the feasibility and effectiveness of our proposed method. This method can construct multi-turn dialogues based on medical, financial, and legal QAs, thereby alleviating the dialogue scarcity in other application domains.
  • To better assess the quality of SmileChat, we collect a real-life counseling dataset with 82 counseling dialogues to build an authentic test set, PsyTest, which contains 4045 utterances spoken by supporters. Both automatic and human evaluations demonstrate that SmileChat improves the performance of the dialogue system in mental health.
  • We release SmileChat, which comprises 55,165 Chinese multi-turn dialogues with an average of 10.4 turns. Additionally, we make our dialogue model, MeChat, and real-life test set, PsyTest, publicly available.
@@ 2.1 Applications of ChatGPT
ChatGPT has proven to be a powerful AI tool for various NLP tasks since its release. Currently, it is being utilized in several domains, such as conversational AI (Alessa and Al-Khalifa, 2023; Köpf et al., 2023; Chen et al., 2023), education (Küchemann et al., 2023; Eshghie and Eshghie, 2023), code programming (Dong et al., 2023; Yetiştiren et al., 2023) and healthcare (Zhao et al., 2023; Yang et al., 2023).
@@
In summary, ChatGPT has already demonstrated its enormous potential as an intelligent pipeline tool that can significantly advance NLP development, despite having only a restricted API available for researchers.
@@ 2.2 Datasets for Mental Health Support
Liu et al. (2021) first define the emotional support conversation task and then, via crowd-sourcing, construct ESConv, an emotional support conversation dataset containing 1053 dialogues with rich support strategies. However, the collection of ESConv requires high cost and time yet leads to a small-scale dialogue dataset. To this end, Zheng et al. (2022) present an approach for augmenting data scale with informative dialogue posts and then constructing AugESC, a model-synthesized dataset with 102k dialogues. The previous two datasets are limited to English. To facilitate the research in Chinese, hence Sun et al. (2021) crawl QA posts in a public mental health support platform compiling PsyQA.
@@ 3 Method
PsyQA [^2] https://www.xinli001.com/qa, a high-quality Chinese dialogue dataset related to mental health support in the form of one question mapping to multiple answers, has gone through a data anonymization process. Our dataset creation pipeline based on PsyQA comprises three main stages: (1) data preprocessing, (2) dialogue generation, and (3) dialogue retention.
3.1 Data Preprocessing
Considering the distinction between QAs in PsyQA and multi-turn dialogues and the context window limitation of 4096 tokens in ChatGPT, we recommend performing data preprocessing for PsyQA. This process involves wording cleaning and length truncation.
@@ 3.1 Data Preprocessing
This work aims to construct a large-scale, diverse, and high-quality multi-turn conversation corpus using the proposed smile method based on PsyQA. While QA can be considered a single-turn conversation between a real help-seeker and a supporter, there are differences in wording compared to actual multi-turn conversations. For instance, the term "楼主" (literally meaning "thread starter") frequently appears in QA but is rarely used in conversation. Therefore, we propose a two-stage process to clean the wording in PsyQA, mitigating linguistic discrepancies before rewriting QA into multi-turn conversations. This process includes both automatic and manual cleaning procedures. For detailed content, see Appendix A.
@@
After conducting a statistical analysis of the PsyQA dataset released by Sun et al. (2021), we find that 757 QAs had a total length exceeding 1800 characters. Additionally, we identify 9 QAs in which the total discourse length exceeds 4000 characters. Furthermore, the model gpt-3.5-turbo [^3] The model we use is gpt-3.5-turbo-0613. has a maximum context length of 4096 tokens. To ensure reliable and smooth rewrites, we limit the length of the QA pairs, maximizing the number of rewritten dialogue turns. Specifically, we truncate the length of the QA pairs at 1800 characters and truncate any excess text. This control ensures that the generated text is limited to approximately 2,000 tokens. It is worth noting that the PsyQA data used in this study undergo a data preprocessing process.
3.2 Dialogue Generation
3.2.1 Task Formulation
Considering the remarkable capabilities of ChatGPT and its cost-effective, plug-and-play characteristics, we consider ChatGPT to be a suitable choice for this task compared to other alternative models. Our study is designed to prompt ChatGPT to rewrite a single-turn dialogue into a multi-turn one. Generally, text generation can be formulated as follows:
$P_{\mathcal{M}_{\theta}}(Y|x)=\prod_{t=1}^{L}P_{\mathcal{M}_{\theta}}(y_{t}|Y_{<t},x)$ (1)
where $\mathcal{M}$ denotes a large language model with parameters $\theta$. $Y$ represents the output of the model generation, while $L$ signifies the variable length of the output. The symbol $y_{t}$ represents the $t$-th token generated by the model, and $x$ denotes the conditional model input.
3.2.2 Prompt Design
@@
To establish some conventions regarding specific terms for prompt design, we define the "help-seeker" and "supporter" as "求助者" and "支持者" in Chinese, respectively. A single-turn dialogue is defined as "求助者:$u^{H}$持者:$u^{S}$", where $u^{H}$ and $u^{S}$ represent the utterances of the help-seeker and supporter, respectively, and $H$ and $S$ refer to the help-seeker and supporter.
In this section, we mainly focus on detailing our methodology. Furthermore, we are building two baseline methods for comparison.
Standard Prompt
As its name suggests, the standard prompt does not contain any single-turn dialogues and instead uses only the simplest prompt to generate multi-turn dialogues. The specifics of the standard prompt are illustrated in Figure 8 in Appendix C. The conditional model input in Equation 1 is $x:=(\mathcal{I})$, where $\mathcal{I}$ represents the standard prompt. We simplify the method name as standard and consider this method as our baseline.
Standard Prompt with a Dialogue Topic
Intuitively, feeding a single, fixed prompt into a large language model often results in the generation of low diversity. To address this issue, we collaborate with three professional counselors, refer to existing literature (Rickwood et al., 2007; Pedrelli et al., 2015), and ultimately compile a comprehensive set of dialogue topics. This set comprises 56 distinct types, each accompanied by its corresponding explanation. For more details, please refer to Appendix K. The conditional model input in Equation 1 is $x:=(\mathcal{I,T})$, where $\mathcal{I}$ represents the dialogue topic chosen in uniform sampling. We simplify the method name to stand.${}_{\mathrm{DT}}$ and adopt this method as our baseline, as illustrated in Figure 8 in Appendix C.
smile Method
Our proposed method, referred to as the smile method, instructs the ChatGPT to rewrite single-turn dialogues into multi-turn ones. Figure 2 depicts the specific prompt template. The conditional model input in Equation 1 is $x:=(\mathcal{I,T,D})$, where $\mathcal{T}$ and $\mathcal{D}$ represent the dialogue topics hidden in the QA and single-turn dialogue, respectively.
3.3 Dialogue Retention
ChatGPT may exhibit potential instability during the dialogue generation process, but this occurs very infrequently. Therefore, we employ an automatic filtering mechanism to exclude dialogues that fail to meet our specified requirements, which encompass dialogue format and dialogue turns as detailed in Appendix B.
@@ 4.1 Experimental Setup
To ensure a fair comparison and prevent duplicate instances of the same question with different answers, we first randomly select 500 non-duplicate questions from the first 5,000 QAs in PsyQA. We then randomly choose one answer to serve as the corresponding question’s response. The data samples obtained are employed as seed dialogues, which are subsequently restructured into multi-turn conversations via ChatGPT. We name this 500 sampled data as PsyQA* to distinguish it from the whole dataset.
@@
On the whole, we present three prompt methods in this paper. For each prompt method, we instruct ChatGPT to generate 500 dialogues to study language transformation and dialogue diversity, thereby validating the feasibility and effectiveness of our proposed smile method, respectively. To enhance the diversity of the generated dialogues, we set ChatGPT’s hyperparameters during text generation to the officially recommended default values, with $\tau$ (temperature) = 1.0 and $p$ (top_p) = 1.0.
@@
[…]
Text Representation
Text representation is used for analyzing language transformation and semantic diversity. A dialogue between a help-seeker and a supporter is represented as $d=\{u_{0}^{H},u_{1}^{S},u_{2}^{H},...,u_{t}^{X},...,u_{n-1}^{H},u_{n}^{S}\}$, where $u_{t}^{X}$ represents the utterance of the $t$-th turn spoken by either the help-seeker or supporter. A string of a dialogue can be denoted as $d_{s}=[u_{0}^{H};u_{1}^{S};u_{2}^{H};...;u_{n-1}^{H};u_{n}^{S}]$, where [;][;][ ; ] denotes the operation of textual concatenation.
To obtain the text embedding of a dialogue, we employ OpenAI’s model text-embedding-ada-002444https://platform.openai.com/docs/guides/embeddings, which accepts a maximum context length of 8191. Each dialogue is first preprocessed into a single string without any speaker tokens and is then mapped to a 1536-dimensional vector. For example, to compute the cosine similarity between two different dialogues, we can obtain

### v3

# Huachuan Qiu1,2 Hongliang He1,2 Shuai Zhang1,2 Anqi Li1,2 Zhenzhong Lan2,3

@@ Abstract
Developing specialized dialogue systems for mental health support requires multi-turn conversation data, which has recently garnered increasing attention. However, gathering and releasing large-scale, real-life multi-turn conversations that could facilitate advancements in mental health support presents challenges in data privacy protection and the time and cost involved in crowdsourcing. To address these challenges, we introduce SMILE, a single-turn to multi-turn inclusive language expansion technique that prompts ChatGPT to rewrite public single-turn dialogues into multi-turn ones. Our work begins by analyzing language transformation and validating the feasibility of our proposed method. We conduct a study on dialogue diversity, including lexical features, semantic features, and dialogue topics, demonstrating the effectiveness of our method. Further, we employ our method to generate a large-scale, lifelike, and diverse dialogue dataset named SMILECHAT, consisting of 55k dialogues. Finally, we utilize the collected corpus to develop a mental health chatbot, MeChat. To better assess the quality of SMILECHAT, we collect a small-scale real-life counseling dataset conducted by data anonymization. Both automatic and human evaluations demonstrate significant improvements in our dialogue system and confirm that SMILECHAT is high-quality. Code, data, and model are publicly available at https://github.com/qiuhuachuan/smile.
@@ 1 Introduction
We all know the importance of mental health, and mental health issues have been a persistent concern for human beings (Kessler et al., 2005). Recent advancements in natural language processing (NLP) technology (Ouyang et al., 2022; Ni et al., 2022) have led to the emergence of neural-based conversational AI in mental health support (Liu et al., 2022; Tu et al., 2022). As an innovative solution for mental health, virtual counselors powered by large language models (LLMs) can effectively address accessibility barriers, such as the high cost of treatment and the shortage of experienced professionals to meet the demand. Furthermore, such dialogue systems provide mental health support as an effective and practical online counseling approach for those in need, safeguarding user privacy and mitigating the stigma that often accompanies help-seeking. However, the lack of publicly available, large-scale, diverse, and high-quality multi-turn chat datasets in the mental health support domain hinders the development of specialized dialogue systems.
Motivation
@@
Indeed, many researchers have been pursuing the goal of building a practical and effective conversational agent for mental health. However, the first step in creating such a system is to have training data. Conversations related to mental health support often contain sensitive information and must be kept confidential to safeguard the privacy of individuals seeking help (Lu et al., 2021). Making these conversations publicly available may discourage individuals from seeking support or negatively impact their personal and professional lives once known to people with whom they are acquainted. To facilitate progress in the NLP community, some researchers have attempted to collect various dialogue corpora (Liu et al., 2021; Sun et al., 2021; Zheng et al., 2022) through crowdsourcing, data crawling, or data augmentation to build a dialogue agent capable of providing emotional and mental health support. How to construct a large-scale, diverse, and high-quality multi-turn chat dataset for mental health without effort motivates us to carry out the work as presented in this paper.
Challenges
@@
To be more specific, crowdsourcing conversations (Liu et al., 2021) for emotional support has limitations due to the high cost and time required to train and manage annotators, as well as the difficulty in mimicking real-life interactions, that is, interlocutors may lack an understanding of the dilemma of living with mental disorders. An alternative is crawling QA (Sun et al., 2021) on a public mental health forum for training psychological support models. However, single-turn conversations may not be sufficient for resolving mental health issues, as multiple interaction exchanges are often needed. Multi-turn conversations, which can better simulate real-world conversations, are more practical for training psychological support models. While the post-triggered machine-augmented method (Zheng et al., 2022) can address the limitations of scale and topic diversity, it does not consider the responses of experienced supporters.
Our Approach
@@
To tackle the abovementioned challenges, we introduce SMILE, single-turn to multi-turn inclusive language expansion via ChatGPT, as shown in Figure 1. Specifically, we instruct ChatGPT to transform publicly available long question-answer pairs (public QAs), which can also be viewed as single-turn dialogues, into multi-turn conversations. With the proposed method, we build a large-scale, diverse, and high-quality multi-turn conversation dataset for mental health support.
@@
  • We first present our method (§$\S$3), including data preparation, task definition, and prompt design that elaborates on the SMILE method and other baseline methods.
  • We then demonstrate the feasibility of the SMILE method through language transformation (§$\S$4), showing that the dialogue constructed by this method is lifelike.
  • Subsequently, we demonstrate the effectiveness of the SMILE method by utilizing three diversity indicators (§$\S$5): lexical features, semantic features, and dialogue topics. Following the validation of feasibility and effectiveness, we leverage the SMILE method to generate a large-scale, lifelike, and diverse multi-turn chat dataset, SMILECHAT.
  • Finally, we propose to train a dialogue system to explore the quality of conversation (§$\S$6) and collect a set of 50 anonymized real-life counseling dialogues for model evaluation.
Our Contributions
We make our data, code, and model publicly available. We believe that our work offers a new perspective on constructing a large-scale, lifelike, diverse, and high-quality multi-turn dialogue dataset for mental health within the research community. Our contributions can be summarized as follows:
  • We introduce SMILE, which provides an easy and novel method for alleviating the scarcity of multi-turn conversations in mental health support.
  • Through the analysis of language transformation and dialogue diversity, we verify the feasibility and effectiveness of our proposed method.
  • To better assess the quality of SMILECHAT, we collect small-scale real-life counseling data with 50 anonymized counseling sessions to build a real-life test set, PsyTest. Automatic and human evaluations on the small-scale real-life test set confirm that our proposed dataset is high-quality.
  • We release SMILECHAT, which consists of 55165 Chinese multi-turn dialogues. Our dialogue model, MeChat, and real-life test set, PsyTest, are publicly available.
We highlight that this method can also construct multi-turn dialogues based on medical, financial, and legal QAs, thereby alleviating the dialogue scarcity in other application domains.
@@ 2.1 Applications of ChatGPT
ChatGPT has proven to be a powerful AI tool for various NLP tasks since its release. Currently, it is being utilized in several domains, such as conversational AI (Alessa and Al-Khalifa, 2023; Köpf et al., 2023; Chen et al., 2023), education (Küchemann et al., 2023; Eshghie and Eshghie, 2023), code programming (Dong et al., 2023; Yetiştiren et al., 2023), and healthcare (Zhao et al., 2023; Yang et al., 2023).
@@
In summary, ChatGPT has already demonstrated its enormous potential as an intelligent pipeline tool that can significantly advance NLP development despite having only a restricted API available for researchers.
@@ 2.2 Datasets for Mental Health Support
Liu et al. (2021) first defined the emotional support conversation task and then, via crowdsourcing, constructed ESConv, an emotional support conversation dataset containing 1053 dialogues with rich support strategies. However, the data collection of ESConv requires high cost and time yet leads to a small-scale dialogue dataset. To this end, Zheng et al. (2022) presented an approach for augmenting data scale with informative dialogue posts and then constructing AugESC, a model-synthesized dataset with 102k dialogues. The previous two datasets are limited to English. To facilitate the research in Chinese, hence Sun et al. (2021) crawled QA posts on a public mental health support platform and made the PsyQA dataset publicly available.
@@ 3 Method
PsyQA [^2] https://www.xinli001.com/qa, an open-source and high-quality Chinese single-turn dialogue dataset focused on mental health support, features one question mapped to multiple answers. Our dataset creation pipeline, based on PsyQA, includes three main stages: (1) data preparation, (2) task definition, and (3) prompt design.
3.1 Data Preparation
Considering the distinction between QA within PsyQA and multi-turn dialogues, along with the context window limitation of 4096 tokens in ChatGPT [^3] The model we use is gpt-3.5-turbo-0613., we propose to perform data preprocessing for PsyQA. This process involves wording cleaning and length truncation.
@@ 3.1 Data Preparation
This work aims to construct a large-scale, lifelike, diverse, and high-quality multi-turn conversation corpus using the proposed SMILE method based on PsyQA. While QA can be considered a single-turn conversation between a real help-seeker [^4] The terms help-seeker and client are interchangeable. and a supporter [^5] The terms supporter and counselor are interchangeable., there are differences in wording compared to actual multi-turn conversations. {CJK*}UTF8gbsnFor instance, the term "楼主" (literally meaning "thread starter") frequently appears in QA but is rarely used in conversation. Therefore, we propose a two-stage process to clean the wording in PsyQA, mitigating linguistic discrepancies before rewriting QA into multi-turn conversations. This process includes both automatic and manual cleaning procedures. Please refer to Appendix A for a detailed process.
@@
After a statistical analysis of the PsyQA dataset, we find that 757 QAs have a total length exceeding 1800 characters. Also, we identify 9 QAs, the total discourse length exceeding 4000 characters. However, the model gpt-3.5-turbo has a maximum context length of 4096 tokens. To ensure reliable and smooth rewrites, we limit the length of the QA pairs, maximizing the number of rewritten dialogue turns. Specifically, we truncate the length of the QA pairs at 1800 Chinese characters and truncate any excess text. This control ensures that the generated text is limited to approximately 2000 tokens. It is worth noting that the PsyQA data used in this study undergo a data preprocessing process.
3.2 Task Definition
Let us first denote the input $x$ as a sequence $\{x_{1},x_{2},...,x_{n}\}$, and the output $y$ as a sequence $\{y_{1},y_{2},...,y_{m}\}$. The generation process of the language model can be expressed as the conditional probability distribution $p(y|x)$, which represents the probability of generating output $y$ given the input $x$. Therefore, text generation via the large language model can be formulated as follows:
$p(y|x)=\prod_{t=1}^{m}p(y_{t}|y_{<t},x)$ (1)
where $y_{t}$ represents the $t$-th token generated by the model. Because $x$ is our main focus in this paper, we will demonstrate the details of the prompt design next.
3.3 Prompt Design
In this section, we mainly focus on describing prompt design. In order to provide a clearer understanding of our method in a more controllable setting and elucidate the superiority of introducing single-turn dialogues as a reference, we first design two baseline prompts for comparison.
3.3.1 standard Prompt
As its name suggests, the standard prompt does not contain any single-turn dialogues or specific dialogue topics and instead uses only the most straightforward prompt to generate multi-turn dialogues. The standard prompt is illustrated in Figure 9 in Appendix C. The input in Equation 1 is $x=I$, where $I$ represents the standard prompt. We simplify the method name as standard and consider this method as our baseline.
3.3.2 standardT Prompt
Intuitively, feeding a single, fixed prompt into a large language model often generates text with low diversity. Therefore, we provide a specific dialogue topic for the standardT prompt. The input in Equation 1 is $x=(I,T)$, where $T$ represents the dialogue topic chosen in uniform sampling in the topic set. We simplify the method name to standardT and adopt this method as our baseline, as illustrated in Figure 10 in Appendix C.
3.3.3 SMILE Prompt
Our paper aims to highlight the superiority of introducing single-turn dialogues when generating dialogues. Our proposed method, the SMILE method, instructs the LLM to rewrite single-turn dialogues into multi-turn ones. Figure 2 depicts the concrete prompt details. The input in Equation 1 is $x=(I,T,D)$, where $T$ and $D$ represent the implicit dialogue topic hidden in the QA and single-turn dialogue, respectively.
@@ 4.1 Experimental Setup
Dialogue Topic Collection
To address the issue of monotonous generation in the standard prompt, we propose to enhance diversity by injecting a dialogue topic into the standardT prompt. Consequently, we have collaborated with three professional counselors, reviewed existing literature (Rickwood et al., 2007; Pedrelli et al., 2015), and ultimately compiled a comprehensive set of dialogue topics. This set includes 60 distinct types of dialogue topics, each with a detailed explanation. For more details, please refer to Appendix G.
@@
To ensure a fair comparison among the three methods we propose and prevent duplicate instances of the same question with different answers, we first randomly select 500 non-repetitive questions from the first 5000 QAs in PsyQA. We then randomly choose one answer to serve as the corresponding response. The data samples obtained are employed as seed dialogues, which are subsequently restructured into multi-turn conversations via ChatGPT. We name this 500 sampled data as PsyQA* to distinguish it from the original dataset.
@@
Overall, we present three prompt methods in this paper. For each prompt method, we instruct ChatGPT to generate 500 dialogues to study language transformation and dialogue diversity, thereby validating the feasibility and effectiveness of our proposed SMILE method, respectively. To enhance the diversity of the generated dialogues, we set hyperparameters during text generation to the officially recommended default values, with temperature = 1.0 and top_p = 1.0.
@@
ChatGPT may exhibit potential instability during the text generation process. Therefore, as detailed in Appendix B, we employ an automatic filtering mechanism to exclude dialogues that fail to meet our specified requirements, which encompass dialogue format and dialogue turns. When a dialogue falls short of these criteria, we prompt ChatGPT to generate the dialogue anew until it aligns with our specified requirements only in the initial 500 dialogues. This measure, therefore, ensures that each method produces an equal number of 500 dialogues for preliminary analysis.
Text Embedding
A multi-turn dialogue between a help-seeker and a supporter is represented as
$d=\{u_{1},r_{1},...,u_{i},r_{i},...,u_{n},r_{n}\}$ (2)
where $u_{i}$ and $r_{i}$ represents the utterances of the $i$-th turn spoken by the help-seeker and supporter, respectively. A string of dialogue without any speaker role tokens can be denoted as $d_{s}=[u_{1};r_{1};u_{2};r_{2};...;u_{n};r_{n}]$, where [;][;][ ; ] denotes the operation of textual concatenation.
To obtain the text embedding of a dialogue, we employ OpenAI’s model text-embedding-ada-002666https://platform.openai.com/docs/guides/embeddings, which accepts a maximum context length of 8191. Each dialogue is first preprocessed into a single string without any speaker tokens and is then mapped to a 1536-dimensional vector. For example, to compute the cosine similarity between two different dialogues, we can obtain